# AGFSync: Leveraging AI-Generated Feedback for Preference Optimization in Text-to-Image Generation

### v1

@@ 3.3 DPO Alignment
$L_{\mathrm{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{w})\sim\mathcal{D},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})}$ (8)
$\log\sigma(-\beta T\omega(\lambda_{k})(\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\theta}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}-\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}$
@@
where $\bm{x}^{*}_{t}=\alpha_{t}\bm{x}^{*}_{0}+\sigma_{t}\bm{\epsilon}^{*}$, $\bm{\epsilon}^{*}\sim\mathcal{N}(0,\bm{I})$. Here, $\alpha_{t}$ and $\sigma_{t}$ are the noise scheduling functions as defined in [23]. Consequently, $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$. Similar to [27], we incorporate $T$ and $\omega(\lambda_{k})$ into the constant $\beta$.
@@ Appendix 0.B Experimental Environments and Settings
Number of Inference Steps
50
Images per Prompt
8
Sampling Precision
FP16
SDXL Resolution
$1024\times 1024$
SD v1.4 & SD v1.5 Resolution
$512\times 512$
SDXL Batch Size
64
SD v1.4 & SD v1.5 Batch Size
128
Max Training Steps
1000
SDXL Learning Rate
0.000001
SD v1.4 & SD v1.5 Learning Rate
0.0000005
Learning Rate Scheduler
Cosine
Mixed Precision
FP16
GPUs for Training
8$\times$×NVIDIA A100(80G)
@@ 0.D.3 Results of Training Models with Different Amounts of Data
Our preference dataset construction method achieved a 100% data conversion efficiency, highlighting the importance of high data conversion efficiency. To further demonstrate this, we randomly selected different proportions of data from our dataset for experiments. Moreover, following the strategy introduced in "DreamSync", we filtered prompts capable of generating high-quality images based on two sets of thresholds (VQA score > 0.85 and Aesthetic score > 0.5, and VQA score > 0.9 and Aesthetic score > 0.6) to construct a preference dataset for training, using SD v1.5 as the base model and conducting the experiment on the TIFA benchmark. The experimental results are shown in Tab. 10. The results indicate that although the Aesthetic scores surpassed the scenario of using all data when trained with 60% and 80% of the data, considering the CLIP score, VQA score, and Aesthetic score together, the larger the volume of data, the better the overall performance of the model. Especially, the improvement in model performance when increasing data usage from 60% to 100% is significantly greater than that from 0% to 60%. Notably, with only our limited constructed preference dataset (i.e. 20% of the preference dataset), we had significant improvements in all aspects, further demonstrating AGFSync’s efficiency.
@@
0 -
27.0
77.1
48.0
20% Random Sample
27.2(+0.2)
77.4(+0.3)
49.1(+1.1)
33.3% $s_{VQA}>0.9,s_{Aes.}>0.6$
27.1(+0.1)
77.4(+0.3)
49.1(+1.1)
40% Random Sample
27.1(+0.1)
77.5(+0.4)
48.8(+0.8)
60% Random Sample
27.1(+0.1)
77.6(+0.5)
49.3 (+1.3)
60.6% $s_{VQA}>0.85,s_{Aes.}>0.5$
27.3 (+0.3)
77.7(+0.6)
47.6(--0.4)
80% Random Sample
27.3 (+0.3)
78.3(+1.2)
49.2(+1.2)
100% -
27.3 (+0.3)
78.7 (+1.6)
49.1(+1.1)

### v3

# Equal contribution    Corresponding author

@@ 3.3 DPO Alignment
$L_{\mathrm{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})}$ (8)
$\log\sigma(-\beta T\omega(\lambda_{t})(\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\theta}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}-\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}$
@@
where $\bm{x}^{*}_{t}=\alpha_{t}\bm{x}^{*}_{0}+\sigma_{t}\bm{\epsilon}^{*}$, $\bm{\epsilon}^{*}\sim\mathcal{N}(0,\bm{I})$. Here, $\alpha_{t}$ and $\sigma_{t}$ are the noise scheduling functions as defined in [23]. Consequently, $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$. Similar to [27], we incorporate $T$ and $\omega(\lambda_{t})$ into the constant $\beta$.
@@ Appendix 0.B Experimental Environments and Settings
Number of Inference Steps 50
Images per Prompt 8
Sampling Precision FP16
SDXL Resolution $1024\times 1024$
SD v1.4 & SD v1.5 Resolution $512\times 512$
SDXL Batch Size 64
SD v1.4 & SD v1.5 Batch Size 128
Max Training Steps 1000
SDXL Learning Rate 0.000001
SD v1.4 & SD v1.5 Learning Rate 0.0000005
Learning Rate Scheduler Cosine
Mixed Precision FP16
GPUs for Training 8$\times$×NVIDIA A100(80G)
@@ 0.D.3 Results of Training Models with Different Amounts of Data
Our preference dataset construction method achieved a 100% data conversion efficiency, highlighting the importance of high data conversion efficiency. To further demonstrate this, we randomly selected different proportions of data from our dataset for experiments. Moreover, following the strategy introduced in "DreamSync", we filtered prompts capable of generating high-quality images based on two sets of thresholds (VQA score > 0.85 and Aesthetic score > 0.5, and VQA score > 0.9 and Aesthetic score > 0.6) to construct a preference dataset for training, using SD v1.5 as the base model and conducting the experiment on the TIFA benchmark. The experimental results are shown in LABEL:tab:different_rate_data_train_result. The results indicate that although the Aesthetic scores surpassed the scenario of using all data when trained with 60% and 80% of the data, considering the CLIP score, VQA score, and Aesthetic score together, the larger the volume of data, the better the overall performance of the model. Especially, the improvement in model performance when increasing data usage from 60% to 100% is significantly greater than that from 0% to 60%. Notably, with only our limited constructed preference dataset (i.e. 20% of the preference dataset), we had significant improvements in all aspects, further demonstrating AGFSync’s efficiency.
@@
0 - 27.0 77.1 48.0
20% Random Sample 27.2(+0.2) 77.4(+0.3) 49.1(+1.1)
33.3% $s_{VQA}>0.9,s_{Aes.}>0.6$ 27.1(+0.1) 77.4(+0.3) 49.1(+1.1)
40% Random Sample 27.1(+0.1) 77.5(+0.4) 48.8(+0.8)
60% Random Sample 27.1(+0.1) 77.6(+0.5) 49.3 (+1.3)
60.6% $s_{VQA}>0.85,s_{Aes.}>0.5$ 27.3 (+0.3) 77.7(+0.6) 47.6(--0.4)
80% Random Sample 27.3 (+0.3) 78.3(+1.2) 49.2(+1.2)
100% - 27.3 (+0.3) 78.7 (+1.6) 49.1(+1.1)